# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 1.7 Final conclusions
Apart from 1.1 with $m=p$ a prime, there are many other evidences for Grothendieck-Katz conjecture. It is verified for “suitable direct factors” of Gauss-Manin connections, see [Kat72], and connections on rank one vector bundles, see [CC85]. Moreover, if the $p$-curvature of a given differential equation is zero for all but a finite number of primes then its singularities are Fuchsian and all the residue matrices are diagonolizable with eigenvalues in $\mathbb{Q}$, see [Kat96, Chapter 9], see also [Kat82, Theorem 8.1]. These are true if the differential equation has algebraic solutions. In other words, the local data of a differential equation with zero $p$-curvatures, looks like a local data of a linear differential equation with only algebraic solutions. In particular, this implies that differential equations which are uniquely determined by their local data (rigid differential equations) satisfy the Grothendieck-Katz conjecture. Therefore, if there is a counterexample to this conjecture we must look for it among non-rigid differential equations, that is, they depend one some auxiliary parameters. Example of such differential equations come from Painlevé VI, Heun and Lamé equations, see [MR10, MR12]. Among these equations, the Lamé equation is the most simple one:
@@
where $p(z)=4z^{3}-g_{2}z-g_{3}$. It depends on the parameters $g_{2},g_{3}\in\mathbb{C},n\in\mathbb{Q},B\in\mathbb{C}$. Our search, with no specific strategy, gave us only Lamé equations with algebraic solutions. These are classified in [BvdW04, Table 4]. We also found the Lamé equation with $B=0,\ n=\frac{7}{4},g_{2}=0,g_{3}=1$ which is not listed in the mentioned reference.  [^7] For further details see Section 2. 1.1 can be reformulated for Gauss-Manin connections, and its converse provides us with a conjectural description of Gauss-Manin connections. This conjecture with $k=1$ appears in [And89, Appendix of Chapter V] and the main evidence for this comes from [Kat72].
@@ 1.8 Appendix: A finitely generated ring
We would like to highlight a very particular case of 1.7 which results in the following elementary statement. Recall that $[a]_{k}:=\frac{a(a-1)\cdots(a-k+1)}{k!}$.
@@
Let $a$ be a positive rational number. The $\mathbb{Z}$-algebra generated by $[a]_{k},\ k\in\mathbb{N}_{0}$ is of the form $\mathbb{Z}[\frac{1}{N}]$, for some $N\in\mathbb{N}$ which is a product of distinct primes. In particular, the number of primes appearing in the denominator of $[a]_{k},\ k\in\mathbb{N}$ is finite.
@@
Let $a=\frac{d}{n},\ {\rm gcd}(n,d)=1$. We first prove that such a $\mathbb{Z}$-algebra $A$ is finitely generated by $[a]_{k},\ k=0,1,2,\ldots,d$ and $\frac{1}{n}$. For $k>d$ we have
@@
which follows from (8) applied to the equality $y^{n}-z^{d}=0$ with $y:=z^{a}$. This implies that $n[a]_{k}$ is a $\mathbb{Z}$-linear combination of products of $[a]_{s},\ s<k$. Therefore, $A$ is finitely generated. For rational numbers $a_{i}=\frac{d_{i}}{n_{i}},\ {\rm gcd}(n_{i},d_{i})=1$ we have $A=\mathbb{Z}[a_{1},a_{2},\ldots]=\mathbb{Z}[\frac{1}{n_{1}},\frac{1}{n_{2}},\cdots]$ which follows from $s_{i}n_{i}+r_{i}d_{i}=1$ or equivalently $s_{i}+r_{i}a_{i}=\frac{1}{n_{i}}$ for some $r_{i},s_{i}\in\mathbb{Z}$. In a similar way, $\mathbb{Z}[\frac{1}{n_{1}},\frac{1}{n_{2}}]=\mathbb{Z}[\frac{1}{[n_{1},n_{2}]}]$ and then by induction $A=\mathbb{Z}[\frac{1}{N}]$ for some $N\in\mathbb{N}$. If $N=N_{1}^{2}N_{2}$ is not square free then $\mathbb{Z}[\frac{1}{N_{1}^{2}N_{2}}]=\mathbb{Z}[\frac{1}{N_{1}N_{2}}]$ and the result follows. ∎
@@
The main interest in the literature is those cases in which $s=n_{1}+n_{2}+\cdots+n_{k}$, see for instance [Zud02].
@@ 1.9 Appendix: Eisenstein theorem in many variables
$\Delta\cdot y_{n}=\hbox{ a polynomial of degree $\leq|n|$ in $y_{i},\ i<n$, with coefficients in ${\mathfrak{R}}$},$
@@
By induction we can show that $y_{n}$ has a pole order at most $12|n|-$ 1 at $\Delta$. For $|n|=1$ this follows from the above equality. If this is true for all $m<n$ then $\Delta y_{n}$ is a sum of monomials $y_{i_{1}}y_{i_{2}}\cdots y_{i_{k}}$ with $i_{1}+i_{2}+\cdots+i_{k}\leq n$ and $i_{1},i_{2},\ldots,i_{k}<n$ and coefficients in ${\mathfrak{R}}$. If $k=1$ then $22|i_{1}|-1\leq 2|n|-$ 2 and we are done. If $k\geq 2$ then $22|i_{1}|-1+2|i_{2}|-1\cdots+2|i_{k}|-1\leq 2|n|-k\leq 2|n|-$ 2 and we are done again. It follows that $N:=\Delta^{2}$ satisfy the desired property. ∎
@@
For $10<r<$ 1 by definition $\langle r\rangle=1$. For an integer $a$ we denote by $\bar{a}$ it unique representative modulo $d$ in $\{0,1,\ldots,d-1\}$. The following is a special case of [Mov21a, Theorem 13.4] with $n=0,\ k=1$
@@
For $\beta\in\mathbb{N}_{0}$ with $20\leq\beta\leq d-$ 2 we have
(16) $\frac{x_{2}^{\beta}(t)}{f^{\prime}(x_{2}(t))}-\frac{x_{1}^{\beta}(t)}{f^{\prime}(x_{1}(t))}={{{\sum}}}_{a\in\mathbb{N}_{0}^{d},\ d\nmid(\beta+1-\sum_{i=1}^{d}ia_{i})}\left(\frac{1}{a_{1}!a_{1}!\cdots a_{d}!}D_{a}{\sf p}_{a}\right)t^{a},$

### v2

@@ 1.7 Final conclusions
Apart from 1.1 with $m=p$ a prime, there are many other evidences for Grothendieck-Katz conjecture. It is verified for “suitable direct factors” of Gauss-Manin connections, see [Katz1972], and connections on rank one vector bundles, see [ChudChud1985-1]. Moreover, if the $p$-curvature of a given differential equation is zero for all but a finite number of primes then its singularities are Fuchsian and all the residue matrices are diagonolizable with eigenvalues in $\mathbb{Q}$, see [Katz1996, Chapter 9], see also [Katz1982, Theorem 8.1]. These are true if the differential equation has algebraic solutions. In other words, the local data of a differential equation with zero $p$-curvatures, looks like a local data of a linear differential equation with only algebraic solutions. The differential equations over $\mathbb{P}^{1}$ which are uniquely determined by their local data (rigid differential equations) satisfy the Grothendieck-Katz conjecture, see [Katz1996, Chapter 9]. Therefore, if there is a counterexample to this conjecture over $\mathbb{P}^{1}$ we must look for it among non-rigid differential equations, that is, they depend one some auxiliary parameters. Example of such differential equations come from Painlevé VI, Heun and Lamé equations, see [ho16, ho17]. Among these equations, the Lamé equation is the most simple one:
@@
where $p(z)=4z^{3}-g_{2}z-g_{3}$. It depends on the parameters $g_{2},g_{3}\in\mathbb{C},n\in\mathbb{Q},B\in\mathbb{C}$. Our search, with no specific strategy, gave us only Lamé equations with algebraic solutions. These are classified in [Beukers2004, Table 4]. We also found the Lamé equation with $B=0,\ n=\frac{7}{4},g_{2}=0,g_{3}=1$ which is not listed in the mentioned reference.  [^5] For further details see Chapter 2. 1.1 can be reformulated for Gauss-Manin connections, and its converse provides us with a conjectural description of Gauss-Manin connections. This conjecture with $k=1$ appears in [Andre1989, Appendix of Chapter V] and the main evidence for this comes from [Katz1972].
@@ 1.8 Appendix: A finitely generated ring
We would like to highlight a very particular case of 1.7 which results in the following elementary statement.
@@
Let $a$ be a positive rational number. The $\mathbb{Z}$-algebra generated by $a]_{k}:=\frac{a(a-1)\cdots(a-k+1)}{k!},\ k\in\mathbb{N}_{0}$ is of the form $\mathbb{Z}[\frac{1}{N}]$, for some $N\in\mathbb{N}$ which is a product of distinct primes. In particular, the number of primes appearing in the denominator of $a]_{k},\ k\in\mathbb{N}$ is finite.
@@
Let $a=\frac{d}{n},\ {\rm gcd}(n,d)=1$. We first prove that such a $\mathbb{Z}$-algebra $A$ is finitely generated by $a]_{k},\ k=0,1,2,\ldots,d$ and $\frac{1}{n}$. For $k>d$ we have
@@
which follows from (1.8) applied to the equality $y^{n}-z^{d}=0$ with $y:=z^{a}$. This implies that $n[a]_{k}$ is a $\mathbb{Z}$-linear combination of products of $a]_{s},\ s<k$. Therefore, $A$ is finitely generated. For rational numbers $a_{i}=\frac{d_{i}}{n_{i}},\ {\rm gcd}(n_{i},d_{i})=1$ we have $A=\mathbb{Z}[a_{1},a_{2},\ldots]=\mathbb{Z}[\frac{1}{n_{1}},\frac{1}{n_{2}},\cdots]$ which follows from $s_{i}n_{i}+r_{i}d_{i}=1$ or equivalently $s_{i}+r_{i}a_{i}=\frac{1}{n_{i}}$ for some $r_{i},s_{i}\in\mathbb{Z}$. In a similar way, $\mathbb{Z}[\frac{1}{n_{1}},\frac{1}{n_{2}}]=\mathbb{Z}[\frac{1}{[n_{1},n_{2}]}]$ and then by induction $A=\mathbb{Z}[\frac{1}{N}]$ for some $N\in\mathbb{N}$. If $N=N_{1}^{2}N_{2}$ is not square free then $\mathbb{Z}[\frac{1}{N_{1}^{2}N_{2}}]=\mathbb{Z}[\frac{1}{N_{1}N_{2}}]$ and the result follows. ∎
@@
The main interest in the literature is those cases in which $s=n_{1}+n_{2}+\cdots+n_{k}$, see for instance [zu02].
The converse of 1.10 must be also true, that is, if the $\mathbb{Z}$-algebra generated by $a]_{k},\ k\in\mathbb{N}_{0}$ is finitely generated then $a$ must be a rational number. For the proof we might use:
Theorem 1.4 (Kronecker).
Let $\mathfrak{k}$} be a number field and $a$ be an element in the ring of integers ${\mathcal{O}}_{\mathfrak{k}}$ of $\mathfrak{k}$}. The number $a$ is in $\mathbb{Z}$ if and only if for all except a finite number maximal prime ideals ${\mathfrak{p}}\subset{\mathcal{O}}_{\mathfrak{k}}$, its class in ${\mathcal{O}}_{\mathfrak{k}}/{\mathfrak{p}}\cong\mathbb{F}_{p^{n}}$ is in $\mathbb{F}_{p}$.
Another useful version of this theorem is:
Theorem 1.5.
Let $P$ be an irreducible non-constant polynomial in $\mathbb{Z}[x]$ of degree $d$. If for all except a finite number primes $p$, $P$ has $d$ roots in $\mathbb{F}_{p}$ then $P$ is of degree $1$, and hence, it has a unique rational root.
1.4 is known under the name Kronecker’s density theorem, see [Esnault2023] in which the author refers to M. Bauer. [^6] M. Bauer, Über einen Satz von Kronecker. Arch. d. Math. u. Phys. 6, 212–222 (1904). For the proof one mainly proves its generalizations such as Frobenius and Chebotarev density theorems. The author was not able to find an elementry proof.
@@ 1.9 Appendix: Eisenstein theorem in many variables
[…]
@@
By induction we can show that $y_{n}$ has a pole order at most 2|n|12|n|-1 at $\Delta$. For $|n|=1$ this follows from the above equality. If this is true for all $m<n$ then $\Delta y_{n}$ is a sum of monomials $y_{i_{1}}y_{i_{2}}\cdots y_{i_{k}}$ with $i_{1}+i_{2}+\cdots+i_{k}\leq n$ and $i_{1},i_{2},\ldots,i_{k}<n$ and coefficients in $\mathfrak{R}$}. If $k=1$ then $2|i_{1}|-1\leq 2|n|-2$ and we are done. If $k\geq 2$ then $2|i_{1}|-1+2|i_{2}|-1\cdots+2|i_{k}|-1\leq 2|n|-k\leq 2|n|-2$ and we are done again. It follows that $N:=\Delta^{2}$ satisfy the desired property. ∎
@@
For $0<r<1$ by definition $\langle r\rangle=1$. For an integer $a$ we denote by $\bar{a}$ it unique representative modulo $d$ in $\{0,1,\ldots,d-1\}$. The following is a special case of [ho13, Theorem 13.4] with n=0,k=1n=0,\ k=1
@@
For $\beta\in\mathbb{N}_{0}$ with $0\leq\beta\leq d-2$ we have
(1.16) $\frac{x_{2}^{\beta}(t)}{f^{\prime}(x_{2}(t))}-\frac{x_{1}^{\beta}(t)}{f^{\prime}(x_{1}(t))}={{{\sum}}}_{a\in\mathbb{N}_{0}^{d},\ d\nmid(\beta+1-\sum_{i=1}^{d}ia_{i})}\left(\frac{1}{a_{1}!a_{1}!\cdots a_{d}!}D_{a}{\sf p}_{a}\right)t^{a},$